# GraphXAIN: Narratives to Explain Graph Neural Networks

### v2

@@ 4.3 Graph Explainer
To obtain subgraph and feature importance explanations, we used GNNExplainer [36], the current state-of-the-art method for explaining GNN models. The GNNExplainer formulates the explanation task as an optimisation problem, aiming to maximise the mutual information between the explanatory subgraph and a subset of node features, relative to the input graph under consideration. The training process for GNNExplainer was conducted for 200 epochs, adhering to the default settings recommended for this explainer by the authors. Nevertheless, in our framework, any graph explainer may be used as long as it provides an explanatory subgraph.
4.4 Large Language Models and Prompts
XAI Narratives are derived from the GNNExplainer output, with the use of LLM (GPT-4o). Using a specifically developed prompt, we guide the LLM to produce coherent and relevant narratives that accurately reflect the graph model’s prediction. The language model is provided with a prompt containing the target node’s feature values, along with feature importance and the subgraph derived from GNNExplainer. Additionally, feature values of the nodes within the subgraph and the final GNN model prediction for the target node are included. In our visualisations, we use the seven most important features and restrict the subgraph to seven nodes, aligning with Miller’s (1956) [24] theory on cognitive limits, which suggests that seven pieces of information represent an optimal amount for receiving, processing and recalling. The full prompt schema is presented in Appendix A.
@@ 5 Results
In the following, we provide examples demonstrating various automatically generated GraphXAINs to explain the GNN model’s final prediction of the target node. It is important to emphasize that the results presented are not selectively chosen. From the graph dataset, five node indices were randomly selected. The GraphXAINs are subsequently generated for each of the selected nodes and are presented in Figure 3 and Figure 4. For the remaining GraphXAIN examples, see Appendix B.1.
@@
The results of the GraphXAINs demonstrate the effectiveness of generating XAI narrative-based explanations from the context-free subgraphs and feature importance values. Unlike other GNN XAI descriptive outputs, which typically present static feature values without context, the outputs generated by our method provide a coherent XAI Narrative. Not only do these narratives articulate which features contributed most to the model’s prediction, but also explain how and why these features combined in a cause-and-effect manner. For instance, the narrative for a high-salary prediction highlights the player’s field and performance statistics, such as position and the number of games played, while crucially contextualising these within broader patterns of team dynamics and social interactions.
@@
Figure 5 presents a possible XAI Description for the GNN model’s final prediction of node 57, outlining key features contributing to the target label prediction. The XAI Description is presented in a static, context-free, and non-coherent format, emphasising the fundamental distinctions between a description and a narrative. The superiority of the XAI Narrative presented at bottom of the Figure 1 over the provided XAI Description is evident. Thus GraphXAIN goes beyond the XAID as shown in Figure 5.
@@
This approach ensures that the explanation is not limited to mere numerical and visual descriptions, but instead offers a comprehensive, story-like narrative that enables the end-user to understand the model’s decision-making process more intuitively. By conveying the reasoning behind the prediction, the narrative XAI explanation helps to bridge the gap between technical model outputs and user comprehension, which is a key purpose of effective narrative communication [4].
@@ 6 Conclusion and Future Work
In this work, we have introduced GraphXAIN, a novel solution that transforms explanatory subgraphs and feature importance metrics into coherent XAI Narrative explanations for GNN models. The proposed framework addresses a significant gap in the current landscape of Explainable AI by moving beyond technical, descriptive outputs and providing intuitive, story-like explanations that enhance user comprehension and trust. Our results demonstrate that XAI Narratives can deliver clearer insights into the decision-making processes of complex GNN models, making these processes more accessible to non-expert users.
For future work, it will be essential to test our framework on other real-world graph datasets to further validate its effectiveness across different domains. Moreover, due to the unstructured nature of graph data, LLMs may be more prone to generate hallucinations, since they struggle to interpret such type of data with the same accuracy as structured inputs [13]. This issue could potentially be mitigated by integrating the proposed method with an approach tailored for graph inputs for LLMs like for instance, the G-Retriever [9]. Additionally, there is a need to develop robust quantitative and qualitative methodologies to assess the accuracy of the generated XAI Narratives.
@@
The data and code used in this research are available at: https://github.com/ADMAntwerp/GraphXAIN
@@ Appendix A Prompt Schema
Figure 6 presents the schema of the prompt used within the proposed framework.
@@ B.1 GraphXAIN
In this section, we present the remaining examples generated by our framework. Figure 7 and Figure 8 provide further insight into the interpretability of the GNN’s model prediction by presenting coherence narratives, which demonstrates the GraphXAIN’s consistency across a range of randomly sampled nodes. Each example was generated following the same methodology, with nodes randomly sampled using the random.sample(range(0, len(data.x)), 5) function and set_seed(42) to ensure reproducibility. The examples showcase the explanatory depth of GraphXAINs in capturing the reasoning behind the GNN model’s predictions.
@@
For example, to obtain a connected subgraph for node 379, GNNExplainer requires at least 14 nodes in this particular example. In this larger subgraph, previously disconnected nodes 7 and 263, become connected to the target node 379, revealing their substantial influence on the final prediction. However, as the subgraph expands, it becomes increasingly complex and challenging to interpret intuitively. Thus, a complementary natural language narrative is needed to provide a coherent understanding of the GNN model’s prediction. Figure 9 presents GraphXAIN’s output for a connected subgraph containing the 14 most influential nodes for the target node’s 379 prediction.
@@
This section presents three examples of textual explanations proposed by Giorgi et al. (2024) [6] for counterfactual explanations of node classification using a GNN model trained on the Cora dataset. The Cora dataset is a standard citation network dataset in which nodes represent research papers, and edges denote citation relationships, with the task being to classify each paper into one of seven categories. Since the authors do not provide any output examples in their article, the examples presented below are taken from the article’s official GitHub repository. These examples were not selectively chosen but represent the first three explanations generated by the best-performing model, as reported in the paper and available in the article’s official GitHub repository222https://github.com/flaat/llm-graph-cf/blob/main/data/results/cf-gnn_Qwen2.5-14B-Instruct-GPTQ-Int4_cora_Response.json:
@@
Again, the explanations presented by He et al. (2024) [10] are delivered in a context-free and fact-based manner, without illustrating underlying relationships or cause-and-effect relationships, thereby categorising them as XAI Descriptions rather than XAI Narratives.
@@
The presented examples align with the proposed XAI Narrative definition. However, the TAGExplainer approach applies only to text-attributed graphs and does not provide an explanatory subgraph or feature importance.

### v3

@@ 4.3 Graph Explainer
To obtain subgraph and feature importance explanations, we use GNNExplainer [44], the current state-of-the-art method for explaining GNN model predictions. The GNNExplainer formulates the explanation task as an optimization problem that maximizes the mutual information between the explanatory subgraph and a subset of node features, relative to the input graph under consideration. In both scenarios (classification and regression), the GNNExplainer training process was conducted for 200 epochs, adhering to the default settings recommended for this explainer by the authors[44]. Different shades of the edges in the subgraph indicate the strength of importance of each connection, with darker edges indicating higher strength. Nevertheless, in our framework, any graph explainer may be used as long as it provides an explanatory subgraph and feature importance values.
4.4 Graphs and Large Language Models
XAI Narratives are derived from the GNNExplainer output, with the use of a LLM (GPT-4o). Using a specifically developed prompt, we guide the LLM in generating the GraphXAIN that accurately reflects the graph model’s prediction. The language model is provided with a prompt containing the target node’s feature values, along with feature importance and the subgraph derived from GNNExplainer.
Additionally, feature values of the nodes within the subgraph and the final GNN model prediction for the target node are included. In our visualisations, we use the seven most important features and restrict the subgraph to seven nodes, aligning with Miller’s (1956) [29] theory on cognitive limits, which suggests that seven pieces of information represent an optimal amount for receiving, processing and recalling. The full prompt schema used to generate GraphXAINs is presented in Appendix A.
@@ 5 Results
In the following, we provide examples presenting various automatically generated GraphXAINs to explain the GNN model’s prediction for the target node. It is important to clarify that the presented results are not selectively chosen - rather, a subset of nodes (subject to explanation) is randomly drawn from both datasets. The GraphXAINs are subsequently generated for each of the selected nodes. Figure 1 and Figure 3 present examples of generated GraphXAINs for node classification and node regression tasks, respectively. For some other GraphXAIN examples, see Appendix B.1.
@@
The results of the GraphXAINs show the effectiveness of generating XAI narrative-based explanations from the context-free subgraphs and feature importance values. Unlike other GNN XAI descriptive outputs, which typically present static feature values, the outputs generated by our method provide a coherent XAI Narrative. Not only do these narratives articulate which features contributed most to the model’s prediction, but also explain how and why these features combined in a cause-and-effect manner. For instance, the narrative for a high-salary prediction highlights the player’s field and performance statistics, such as position and the number of games played, while crucially contextualising these within broader patterns of team dynamics and social interactions.
@@
Figure 4 presents a possible XAI Description for the GNN model’s final prediction of node 57, outlining key features contributing to the target label prediction. The XAI Description is presented in a static, context-free, and non-coherent format, emphasising the fundamental distinctions between a description and a narrative. The advantage of the XAI Narrative presented at bottom of the Figure 1 over the provided XAI Description is evident. Thus GraphXAIN goes beyond the XAI Description as shown in Figure 4.
| Dimension | Question |
| Understandability | Q1: I find the explanation understandable. |
| Trustworthiness | Q2: This explanation increased my trust in the model’s prediction. |
| Insightfulness | Q3: The explanation helped me gain insight into the factors influencing the classification/regression. |
| Satisfaction | Q4: I am satisfied with the explanation’s clarity and thoroughness. |
| Confidence | Q5: I feel confident in the classification/regression after reviewing the explanation. |
| Convincingness | Q6: I find this explanation convincing in justifying why node X is classified as Y/predicted to have a regression score of Y. |
| Communicability | Q7: I find this explanation method suitable for communicating the model’s predictions to others. |
| Usability | Q8: If I work with Graph Neural Network models in the future, I am likely to use this explanation method to explain the model’s prediction. |
@@
This approach ensures that the explanation is not limited to mere numerical and visual descriptions, but instead offers a comprehensive, story-like narrative that enables the practitioners to understand the model’s decision-making process more intuitively. By conveying the reasoning behind the prediction, the XAI Narrative explanation addresses a primary goal in XAI by bridging the gap between technical model outputs and user comprehension, thereby reflecting both the principles of narrative communication and the general objectives of explainability research [6].
Note that in the IMDB dataset, movie titles are available and directly mapped to nodes, whereas in the NBA dataset, player names are hashed, preventing their direct mapping. We hypothesize that having actual names would allow GraphXAIN to produce more effective explanations by drawing on specific player information.
6 Evaluation with human
subjects: a user study
To evaluate the proposed GraphXAIN method, we conducted a user study similar to Baniecki et al. (2023) and Martens et al. (2024)[2, 26]. Twenty active researchers and practitioners from the field of machine learning (ML) and data science (DS) participated in the evaluation survey. Seventy-five percent of the participants are academic ML researchers, remaining 25% are industry ML/DS practitioners. Regarding professional experience, 30% of the participants had between 0 and 2 years of experience, 50% had 3 to 5 years, 15% had 6 to 10 years, and 5% had more than 10 years of experience in the domain.
The purpose of the survey was to evaluate the impact of the proposed GraphXAIN method in comparison with the current state-of-the-art XAI method[44] to explain GNN predictions. Six randomized GNN explanation examples were shown to each participant sequentially. Immediately after viewing each XAI method, participants are asked to rate their agreement with eight questions on a five-point Likert scale, ranging from “Strongly disagree” (1) to “Strongly agree” (5), where higher scores indicate stronger agreement or preference. Each participant was presented with three examples drawn from both classification and regression tasks: 1) a current state-of-the-art GNNExplainer [44] explanation comprising a subgraph and feature-importance attributions, 2) GraphXAIN, which provides a narrative explanation of the model’s prediction, 3) a combined approach that integrates GNNExplainer and GraphXAIN. The order of presenting the three aforementioned methods to the participants was randomized.
Respondents evaluated XAI methods across eight dimensions: understandability, trustworthiness, insightfulness, satisfaction, confidence, convincingness, communicability, and usability. Table 1 presents the survey questions used to measure each dimension. The following section presents the theoretical principles for selecting these evaluation dimensions.
Trustworthiness - Hoffman et al. (2018) note that fostering appropriate trust is one of the core objectives of explanations, underscoring the need to assess whether an explanation effectively increases trust [16].
Insightfulness - Riberio et al. (2016) declare that a good explanation should provide additional insights into model predictions [37].
Satisfaction - Miller (2019) investigates the cognitive and social dimensions that render explanations meaningful to humans, positing that, among others, practitioners and user satisfaction is fundamental to effective explanatory strategies [30].
Confidence - Ribeiro et al. (2024) argue that explanations should enable individuals to gain confidence in the model [36].
Convincingness - Miller (2019) underscores that an explanation must be persuasive to the user [30].
| Dimension | Method Comparison ( vs ) | ) | -value | ||
| Understandability | SubG+FI vs GraphXAIN | 0.004* | |||
| SubG+FI vs SubG+FI+GraphXAIN | 0.003* | ||||
| GraphXAIN vs SubG+FI+GraphXAIN | 1.000 | ||||
| Trustworthiness | SubG+FI vs GraphXAIN | 0.254 | |||
| SubG+FI vs SubG+FI+GraphXAIN | 0.005* | ||||
| GraphXAIN vs SubG+FI+GraphXAIN | 0.072 | ||||
| Insightfulness | SubG+FI vs GraphXAIN | 0.178 | |||
| SubG+FI vs SubG+FI+GraphXAIN | 0.018* | ||||
| GraphXAIN vs SubG+FI+GraphXAIN | 0.251 | ||||
| Satisfaction | SubG+FI vs GraphXAIN | 0.045* | |||
| SubG+FI vs SubG+FI+GraphXAIN | 0.003* | ||||
| GraphXAIN vs SubG+FI+GraphXAIN | 0.586 | ||||
| Confidence | SubG+FI vs GraphXAIN | 0.085 | |||
| SubG+FI vs SubG+FI+GraphXAIN | 0.015* | ||||
| GraphXAIN vs SubG+FI+GraphXAIN | 0.786 | ||||
| Convincingness | SubG+FI vs GraphXAIN | 0.030* | |||
| SubG+FI vs SubG+FI+GraphXAIN | 0.021* | ||||
| GraphXAIN vs SubG+FI+GraphXAIN | 0.314 | ||||
| Communicability | SubG+FI vs GraphXAIN | 0.020* | |||
| SubG+FI vs SubG+FI+GraphXAIN | 0.001* | ||||
| GraphXAIN vs SubG+FI+GraphXAIN | 0.019* | ||||
| Usability | SubG+FI vs GraphXAIN | 0.284 | |||
| SubG+FI vs SubG+FI+GraphXAIN | 0.008* | ||||
| GraphXAIN vs SubG+FI+GraphXAIN | 0.541 |
Communicability - Miller (2019) argues that the information exchanged between the explainer and the explainee should align with the general rules of cooperative conversation, proposed by Grice’s (1975) [11], ensuring that it remains relevant to the explainee’s context and builds upon their prior knowledge [30].
Usability - Miller (2017) argues that individuals evaluate explanations according to pragmatic influences of causality, which encompass criteria such as usefulness [31].
Table 2 presents the survey results along with the statistical significance from Wilcoxon signed-rank tests assessing the differences between XAI methods, conducted at the significance level.
When comparing preferences between GNNExplainer (presented in Table 2 as “SubG+FI”) and GraphXAIN, the latter is preferred in four dimensions: understandability, satisfaction, convincingness, and communicability, while no statistically significant differences are observed in the other four dimensions.
In another comparison, when evaluating the GraphXAIN alone against a combined explanation that integrates GNNExplainer’s outputs with the GraphXAIN narrative (SubG+FI+GraphXAIN), only one dimension, communicability, shows a statistically significant difference, with the combined method being preferred. For the remaining seven dimensions, no significant differences occur.
Moreover, across all eight investigated dimensions, the combined method (SubG+FI+GraphXAIN) is preferred over the GNNExplainer method (SubG+FI). This finding suggests that incorporating GraphXAIN’s narrative component into the technical subgraph and feature importance explanations consistently enhances the overall quality of the explanation.
Furthermore, 95% of participants (19/20) answered “Yes” to the question, “Do you think that the narratives are a useful addition to explaining the GNN model’s predictions?”. This response aligns with the observed advantages of the combined method (SubG+FI+GraphXAIN) over the standalone GNNExplainer outputs.
Additionally, at the significance level, no statistically significant differences are identified between responses for the NBA and IMDb datasets across any condition or question. Overall, these findings indicate that participants’ ratings do not differ meaningfully between the two datasets, which represent distinct graph predictive tasks.
Although the sample size is modest, the survey findings demonstrate that GraphXAIN consistently outperforms the current state-of-the-art graph explanation method, serving as a valuable enhancement to improve understanding, trust, insightfulness, satisfaction, confidence, convincingness, communicability, and usability.
@@ 7 Conclusion and Future Work
In this work, we introduced GraphXAIN, a novel solution that transforms explanatory subgraphs and feature importance metrics into coherent narrative explanations for Graph Neural Network (GNN) models.
GraphXAIN is a model-agnostic and explainer-agnostic approach that complements traditional graph explainer outputs by leveraging Large Language Models (LLMs). Results obtained from generating narrative explanations for both node classification and node regression tasks on real-world datasets demonstrate that these narratives provide clearer insights into the decision-making processes of complex GNN models, making them more accessible to non-expert users. In doing so, our method addresses a significant gap in the current Explainable AI landscape by moving beyond technical, descriptive outputs and offering intuitive, story-like explanations that enhance practitioners’ comprehension and trust.
A survey conducted among machine learning researchers and practitioners highlights the advantages of GraphXAIN over GNNExplainer [44]. At the level, the GraphXAIN method significantly improved understandability, satisfaction, convincingness, and suitability for communicating the model’s predictions. Furthermore, when GNNExplainer’s outputs are combined with GraphXAIN narratives, all eight measured dimensions (understandability, trustworthiness, insightfulness, satisfaction, confidence, convincingness, communicability, and future usability) show significant improvement compared to using GNNExplainer outputs alone, indicating that integration of the XAI Narratives with subgraph and feature importance outputs enhances users’ perceptions of the explanation. Notably, 95% of participants indicate that GraphXAIN is a valuable addition to the explanation of GNN predictions.
In future work, employing quantitative metrics will be essential for a more thorough assessment of the generated narrative explanations [19]. One promising direction is to tailor our method to generate graph counterfactual narratives that describe the minimal change required to alter the classifier’s decision, similar to the approach proposed by Martens et al. (2024) [26]. Moreover, due to the unstructured nature of graph data, LLMs may be more prone to generating hallucinations since they struggle to interpret unstructured inputs as accurately as structured data [18]. This challenge can be mitigated by integrating our approach with methods specifically designed for graph inputs, such as G-Retriever [13].
@@
The data and code used in this research, as well as the survey results, are available at:
https://github.com/ADMAntwerp/GraphXAIN
@@ B.1 GraphXAIN
In this section, we present more examples generated by the GraphXAIN method. Figure 7 and Figure 8 provide further insight into the interpretability of the GNN’s model prediction by presenting coherence narratives, which demonstrates the GraphXAIN’s consistency across a range of randomly sampled nodes for both node classification and node regression tasks. Each example was generated following the same methodology, with nodes randomly sampled using the random.sample(range(0, len(data.x)), 5) function and set_seed(42). The examples showcase the explanatory depth of GraphXAINs in capturing the reasoning behind the GNN model’s predictions.
@@
For example, to obtain a connected subgraph for node 379, GNNExplainer requires at least 14 nodes in this particular example. In this larger subgraph, previously disconnected nodes 7 and 263 become connected to the target node 379, revealing their substantial influence on the final prediction. However, as the subgraph expands, it becomes increasingly complex and challenging to interpret intuitively. Thus, a complementary natural language narrative is needed to provide a coherent understanding of the GNN model’s prediction. Figure 9 presents GraphXAIN’s output for a connected subgraph containing the 14 most influential nodes for the target node’s 379 prediction.
@@
This section presents three examples of textual explanations proposed by Giorgi et al. (2024) [9] for counterfactual explanations of node classification using a GNN model trained on the Cora dataset. The Cora dataset is a standard citation network dataset in which nodes represent research papers, and edges denote citation relationships, with the task being to classify each paper into one of seven categories. Since the authors do not provide any output examples in their article, the examples presented below are taken from the article’s official GitHub repository. These examples are not selectively chosen but represent the first three explanations generated by the best-performing model, as reported in the paper and available in the article’s official GitHub repository333https://github.com/flaat/llm-graph-cf/blob/main/data/results/cf-gnn_Qwen2.5-14B-Instruct-GPTQ-Int4_cora_Response.json:
@@
Again, the explanations presented by He et al. (2024) [14] are delivered in a context-free and fact-based manner without illustrating underlying relationships or cause-and-effect relationships, thereby categorising them as XAI Descriptions rather than XAI Narratives.
@@
The presented examples align with the proposed XAI Narrative definition. However, the TAGExplainer approach applies to text-attributed graphs only and does not provide an explanatory subgraph or feature importance.